Please help solve this equation.
9m^2-2m=0

Answer by Earlsdon(6294)

Try this:
9m^2 - 2m = 0 Factor the m.
m(9m - 2) = 0 Apply the zero products principle.
m = 0 or,
9m - 2 = 0
9m = 2
m = 2/9